# Fantastic LLMs for Preference Data Annotation and How to (not) Find Them

### v1

@@ Abstract
Preference tuning of large language models (LLMs) relies on high-quality human preference data, which is often expensive and time-consuming to gather. While existing methods can use trained reward models or proprietary model as judges for preference annotation, they have notable drawbacks: training reward models remain dependent on initial human data, and using proprietary model imposes license restrictions that inhibits commercial usage. In this paper, we introduce customized density ratio (CDR) that leverages open-source LLMs for data annotation, offering an accessible and effective solution. Our approach uses the log-density ratio between a well-aligned LLM and a less aligned LLM as a reward signal. We explores 221 different LLMs pairs and empirically demonstrate that increasing the performance gap between paired LLMs correlates with better reward generalization. Furthermore, we show that tailoring the density ratio reward function with specific criteria and preference exemplars enhances performance across domains and within target areas.
In our experiment using density ratio from a pair of Mistral-7B models, CDR achieves a RewardBench score of 82.6, outperforming the best trained reward functions from same model class and demonstrating competitive performance against SoTA models in Safety (91.0) and Reasoning (88.0) domains. We use CDR to annotate an on-policy preference dataset with which we preference tune Llama-3-8B-Instruct with SimPO. The final model achieves a 37.4% ($+$15.1%) win rate on ArenaHard and a 40.7% ($+$17.8%) win rate on Length-Controlled AlpacaEval 2.0, along with a score of 8.0 on MT-Bench.
@@ 1 Introduction
This paper introduces customized density ratio (CDR) that leverages the density ratio between off-the-shelf open-source LLMs to efficiently facilitate preference data annotation. Our method uses the log-density ratio between a well-aligned model and less-aligned model to annotate preference data. We show that a higher alignment gap between model pairs yields improved preference signal. This observation, referred to as the “Strong-over-Weak Hypothesis”, is supported by our experiments across 221 model combinations (Figure 1). Notably, log-density ratios between a post-DPO model and a pre-DPO model, known as the DPO implicit reward (Rafailov et al., 2023), have not gained widespread adoption due to limitations in generalizability and high reward variance in reward performance across different model choices (Lambert et al., 2024; Lin et al., 2024). We show that the implicit DPO reward is an empirically suboptimal special case of the strong-over-weak density ratio reward, with its performance variance reducible by consistently choosing a weaker model as the reference (Figure 2). Our findings highlight the importance of selecting model pairs with a sufficient alignment gap and demonstrate flexibility in model selection (SFT, RRHF, SLiC-HF, ORPO, SimPO, KTO, IPO, etc.) for constructing the density ratio reward.
@@
  • Choosing Models via Strong-over-Weak Hypothesis. We propose a general framework leveraging the density ratio between a well-aligned LLM and a less-aligned LLM as a reward signal for annotating preference data. We introduce the “Strong-over-Weak Hypothesis”, which suggests increasing the preference gap between the two LLMs to improve the accuracy of the density ratio reward function. Through extensive experimentation on 221 model pairs, we empirically validate this hypothesis. Our findings can reduce the reward variance seen in existing density ratio methods (e.g., DPO Implicit reward) and offer guidance on selecting effective model pairs for density ratio-based reward functions.
@@ 4.2 Customized Strong-Over-Weak Density Ratio
density ratio 92.2 60.5 82.4 73.8 77.2
density ratio (base) 89.9 65.6 62.8 71.9 71.9
density ratio (sft, base) 79.6 65.6 52.8 70.0 67.0
@@
Table 1: Performance on Reward Bench across multiple dimensions (Chat, ChatHard, Safety, and Reasoning). The overall score is the average of these four. GPT-4 and Claude-3.5 are proprietary models serving as examples of LLM-as-a-judge reward functions. Unless otherwise specified, all models below ArmoRM use Nous-Hermes-2-Mistral-7B-DPO as the strong model and OpenHermes-2.5-Mistral-7B as the weak model for density ratio reward. “base” indicates density ratio between DPO model and base (Mistral-7B-v0.1); “sft, base” indicates ratio between the SFT model and base. CDR (Customized Density Ratio) applies domain-specific guidance (e.g., safety or code/math or chat-hard). Adaptive routing configurations include an "oracle" (ideal routing) and a real-world "router" based on a zero-shot prompted LLM.
@@ 4.3 Alignment with density ratio annotated data
density ratio (sft, base) 23.3 21.3 1720 23.5 564 8.3
CDR (adaptive, sft, base) 27.5 26.7 1888 30.4 607 8.3
density ratio 39.9 40.1 2008 34.6 571 8.1
@@
Table 2: Alignment performance after SimPO training on the Llama-3-Instruct (8B) model. Reward function is used to annotate the online preference dataset, obtained through Best-of-32 sampling. The first row is the baseline performance of Llama-3-Instruct (8B) model using a SoTA trained reward function. CDR (Customized Density Ratio) and density ratio uses NousResearch/Nous-Hermes-2-Mistral-7B-DPO as the strong model and teknium/OpenHermes-2.5-Mistral-7B as the weak model, except when ”sft,base” is indicated which uses ratio between SFT model and base model. CDR applies domain-specific guidance (e.g., safety or code/math). Adaptive routing assigns customized instruction of relevant domain for each example.
@@ A.1 Evaluation
We use ArenaHard score as proxy for a model’s human preferred level, it is shown to have the highest correlation and separability against gold human judgments in ChatArena. While it doesn’t not score individual dimensions of preference, it provides an aggregate signal for overall human preference. The delta is calculated as the difference between strong model and weak model’s arena hard score.
@@
MT-Bench is a multi-turn benchmark that measures model performance on 8 dimensions compared to a reference ground-truth.

### v2

# CDR: Customizable Density Ratios of Strong-over-weak LLMs for Preference Annotation

@@ Abstract
Preference tuning of large language models (LLMs) relies on high-quality human preference data, which is often expensive and time-consuming to gather. While existing methods can use trained reward models or proprietary model as judges for preference annotation, they have notable drawbacks: training reward models remain dependent on initial human data, and using proprietary model imposes license restrictions that inhibits commercial usage. In this paper, we introduce customized density ratio (CDR), a training-free and highly effective method that leverages off-the-shelf LLMs for preference data annotation. Our approach uses the log-density ratio between a better-aligned LLM and a less aligned LLM as a reward signal. We explores 221 different LLMs pairs and empirically demonstrate that increasing the performance gap between paired LLMs correlates with better reward generalization. Furthermore, we show that tailoring the density ratio reward function with specific criteria and preference exemplars enhances performance across domains and within target areas.
In our experiment using density ratio from a pair of Mistral-7B models, CDR achieves a RewardBench score of 82.6, outperforming the best trained reward functions from same model class and demonstrating competitive performance against SoTA models in Safety (91.0) and Reasoning (88.0) domains. We use CDR to annotate an on-policy preference dataset with which we preference tune Llama-3-8B-Instruct with SimPO. Using reward signals from two relatively weak models, our approach pushes Llama-3-8B to achieve a 37.4% ($+$15.1%) win rate on ArenaHard and a 40.7% ($+$17.8%) win rate on Length-Controlled AlpacaEval 2.0, along with a score of 8.0 on MT-Bench.
@@ 1 Introduction
This paper introduces customized density ratio (CDR) that leverages the density ratio between off-the-shelf open-source LLMs to efficiently facilitate preference data annotation. Our method uses the log-density ratio between a better-aligned model and less-aligned model to annotate preference data. We show that a higher alignment gap between model pairs yields improved preference signal. This observation, referred to as the “Strong-over-Weak Hypothesis”, is supported by our experiments across 221 model combinations (Figure 1). Notably, log-density ratios between a post-DPO model and a pre-DPO model, known as the DPO implicit reward (Rafailov et al., 2023), have not gained widespread adoption due to limitations in generalizability and high reward variance in reward performance across different model choices (Lambert et al., 2024; Lin et al., 2024). We show that the implicit DPO reward is an empirically suboptimal special case of the strong-over-weak density ratio reward, with its performance variance reducible by consistently choosing a weaker model as the reference (Figure 2). Our findings highlight the importance of selecting model pairs with a sufficient alignment gap and demonstrate the flexibility to construct the density ratio reward using models trained with various objectives (such as SFT, RRHF, SLiC-HF, ORPO, SimPO, KTO, IPO, etc.)
@@
  • Choosing Models via Strong-over-Weak Hypothesis. We propose a training-free framework to leverage the density ratio between a better-aligned LLM and a less-aligned LLM as a reward signal for annotating preference data. We introduce the “Strong-over-Weak Hypothesis”, which suggests increasing the preference gap between the two LLMs to improve the accuracy of the density ratio reward function. Through extensive experimentation on 221 model pairs, we empirically validate this hypothesis. Our findings can reduce the reward variance seen in existing density ratio methods (e.g., DPO Implicit reward) and offer guidance on selecting effective model pairs for density ratio-based reward functions.
@@ 4.2 Customized Strong-Over-Weak Density Ratio
density ratio (dpo vs. sft) 92.2 60.5 82.4 73.8 77.2
density ratio (dpo vs. base) 89.9 65.6 62.8 71.9 71.9
density ratio (sft vs. base) 79.6 65.6 52.8 70.0 67.0
DPO vs SFT
@@
Table 1: Performance on Reward Bench across multiple dimensions (Chat, ChatHard, Safety, and Reasoning). The overall score is the average of these four. RM-Mistral-7B is the strongest in-class trained reward model initialized from mistralai/Mistral-7B-Instruct-v0.2. ArmoRM-Llama-3-8B is a SoTA reward model scoring second on RewardBench by time of writing. GPT-4 and Claude-3.5 are proprietary models serving as examples of LLM-as-a-judge reward functions. To construct the density ratio, we can use a DPO model (Nous-Hermes-2-Mistral-7B-DPO), an SFT model ( OpenHermes-2.5-Mistral-7B), or a Base model (Mistral-7B-v0.1). We denote specific pairings in the format (dpo vs. sft), which, for example, indicates the density ratio between DPO and SFT models. CDR (Customized Density Ratio) applies domain-specific instructions (e.g., safety or code/math or chat-hard) when taking density ratio. Adaptive routing configurations include an "oracle" (ideal routing) and a real-world "router" based on a zero-shot prompted LLM.
@@ 4.3 Alignment with density ratio annotated data
SFT vs Base
vanilla density ratio 23.3 21.3 1720 23.5 564 8.3
CDR (adaptive) 27.5 26.7 1888 30.4 607 8.3
DPO vs SFT
vanilla density ratio 39.9 40.1 2008 34.6 571 8.1
@@
Table 2: Alignment performance after SimPO training on the Llama-3-Instruct (8B) model. Reward function is used to annotate the online preference dataset, obtained through Best-of-32 sampling. The first row is the performance of the starting model Llama-3-Instruct (8B) model. The second row is the alignment performance of aligning using a SoTA trained reward function. DPO model indicated is NousResearch/Nous-Hermes-2-Mistral-7B-DPO; SFT model is teknium/OpenHermes-2.5-Mistral-7B; Base model is mistralai/Mistral-7B-v0.1. CDR (Customized Density Ratio) applies domain-specific guidance (e.g., safety or code/math) to the vanilla density ratio reward. Adaptive indicates using a routing system to assign instruction from relevant domain for each example.
@@ A.1 Evaluation
We use ArenaHard (Li et al., 2024) score as proxy for a model’s human preferred level, it is shown to have the highest correlation and separability against gold human judgments in ChatArena. While it doesn’t not score individual dimensions of preference, it provides an aggregate signal for overall human preference. The delta is calculated as the difference between strong model and weak model’s arena hard score.
@@
MT-Bench (Zheng2023JudgingLW) is a multi-turn benchmark that measures model performance on 8 dimensions compared to a reference ground-truth.